Publications: C.N. Friedman

Articles Perturbations of the Schrödinger equation by potentials with small support, J. Functional Analysis,

(3) 10(1972), 346-360. Semigroup product formulas, compressions, and continual observations in Quantum mechanics,

Indiana Univ. Math. J., (11) 21(1972), 1001-1011. Renormalized oscillator equations, J. Mathematical Phys., (10) 14(1973), 1378-1380. Continual measurements in the space-time formulation of non-relativistic Quantum mechanics,

Ann. Physics, 98(1976), 87-97. (with John D. Dollard) Product integrals and the Schrödinger equations, J. Mathematical Phys.,

18(1977), 1598-1607. (with John D. Dollard) On strong product integration, J. Functional Analysis, (3) 28(1978), 309-

354. (with John D. Dollard) Product integrals II: contour integrals, J. Functional Analysis, (3) 28(1978),

355-368. (with John D. Dollard) Asymptotic behavior of solutions of linear ordinary differential equations,

J. Math. Anal. Appl., 66(1978), 394-398. (with John D. Dollard) Existence of the Möller wave operators for V (r) = C sin(r^{α})/r^{β}, Ann.

Physics (1) III(1978), 251-266. (with John D. Dollard) Product integration of measures and applications, J. Differential Equations,

31(1979), 418-464. Product Integration with Applications to Differential Equations, Encyclopedia of Applied Mathe-

matics, 10(1979). (with John D. Dollard) A new proof of the invariance principle in scattering theory, Ann. Physics,

119(1979), 71-75. (with John D. Dollard) Scattering of a beam of particles by a potential, J. Mathematical Phys. (8)

20(1979), 1727-1730. (with John D. Dollard) Existence and completeness of the Miller wave operators for radial potentials

satisfying

∫ r|V (r)| dr + ∫ |V (r)| dr < ∞ ,

J. Mathematical Phys. (6) 21(1980), 1336-1339. Asymptotic forms of solutions of certain linear ordinary differential equations, J. Math. Anal. Appl.

(2) 80(1980), 461-476. Second order linear O.D.E. and Riccati equations, J. Math. Anal. Appl. (2) 81(1981), 291-296. Product integration and solution of ordinary differential equations, J. Math. Anal. Appl. (2)

102(1984), 509-518. Existence and approximation of solutions of a class of functional equations including ordinary

differential equations, J. Math. Anal. Appl. (2) 112(1985), 446-454. Some remarks on Pauli data in quantum mechanics, J. Austral. Math. Soc., Ser. B. 39(1989),

298-303. Detecting Hamiltonian Cycles, Applied Mathematics and Computation, 47(1992), 137-143. Sums of Divisors and Egyptian Fractions, J. Number Theory, 44 (1993), 1-12. The Frequency Interpretation in Probability, Advances in Applied Mathematics, 23 (1999), 234-254.

October 23, 2005